Apples are 92 cents a pound, and pears are $1.22 a pound. If I spend $ 4 and the weight of the apples I buy is twice the weight of the pears, how many pounds of pears do I buy?
Round your answer to three decimal places.

Answer by TimothyLamb(4379)

x = pounds of apples
y = pounds of pears
---
x = 2y
0.92x + 1.22y = 4
---
put the system of linear equations into standard form
---
x - 2y = 0
0.92x + 1.22y = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = 2.6143791
y = 1.3071895
---
now, solve by substitution:
---
x = 2y
0.92x + 1.22y = 4
0.92*(2y) + 1.22y = 4
(0.92*2 + 1.22)y = 4
y = 4/(0.92*2 + 1.22)
y = 1.30718954248
x = 2y
x = 2*1.30718954248
x = 2.61437908496
---
solution by substitution:
x = 2.6143791
y = 1.3071895
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---